OpenStudy (anonymous):

4√6/√30 I know it equals 4√5/5 but I don't know the steps. Can someone help me?

OpenStudy (ajprincess):

\[\frac{ 4\sqrt{6} }{ \sqrt{30} }\] \[=\frac{ 4\sqrt{6} }{ \sqrt{6}*\sqrt{5}}\] Do u see anything common in the denominator and numerator? if so what remains when u cancel that?

OpenStudy (anonymous):

5 remains so the denominator would 5 and you would get \[\frac{4\sqrt{5} }{ 5 }\]

OpenStudy (ajprincess):

well actually this is what happens  Is it clear?